OpenStudy (anonymous):

If F(x) = 5x + 4, which of the following is the inverse of F(x)? A. F-1(x) = 4 – 5x B. F-1(x) = x-4/5 C. F-1(x) = 5x – 4 D. F-1(x) = x-5/4

OpenStudy (anonymous):

\[f(x)=5x+4\]\[Let~y=5x+4\]\[y-4=5x\]\[\frac{ y-4 }{ 5 }=x\]\[f^{-1}(x)=\frac{ y-4 }{ 5 }\]\[Therefore,f^{-1}(x)=\frac{ x-4 }{ 5 }\]

OpenStudy (anonymous):

which option do u think is correct? @ShyannaD

OpenStudy (anonymous):

A

OpenStudy (anonymous):

B

OpenStudy (anonymous):

yes :)